# Learning Reward Machines from Partially Observed Optimal Policies

### v1

@@ Abstract
Inverse reinforcement learning is the problem of inferring a reward function from an optimal policy. In this work, it is assumed that the reward is expressed as a reward machine whose transitions depend on atomic propositions associated with the state of a Markov Decision Process (MDP). Our goal is to identify the true reward machine using finite information. To this end, we first introduce the notion of a prefix tree policy which associates a distribution of actions to each state of the MDP and each attainable finite sequence of atomic propositions. Then, we characterize an equivalence class of reward machines that can be identified given the prefix tree policy. Finally, we propose a SAT-based algorithm that uses information extracted from the prefix tree policy to solve for a reward machine. It is proved that if the prefix tree policy is known up to a sufficient (but finite) depth, our algorithm recovers the exact reward machine up to the equivalence class. This sufficient depth is derived as a function of the number of MDP states and (an upper bound on) the number of states of the reward machine. Several examples are used to demonstrate the effectiveness of the approach.
I Introduction
@@
Several frameworks exist for solving complex multi-staged tasks, including hierarchical reinforcement learning (HRL) [1], reward machines (RMs) [2] and linear temporal logic (LTL) specifications [3, 4]. HRL leverages a decomposition of tasks into subtasks, enabling agents to focus on solving smaller, manageable problems before integrating solutions into a higher-level policy [5]. On the other hand, RM and its generalizations [6] encode task-specific knowledge as finite-state machines, capturing temporal dependencies and logical constraints in a concise and interpretable manner, similar to LTL. This structure simplifies policy learning and improves efficiency, especially in environments with long horizons or sparse rewards.
@@
As an extension to inverse reinforcement learning (IRL) [7], one could ask the question of learning RMs, which enables agents to autonomously extract structured representations of tasks, significantly enhancing their ability to solve complex, temporally extended problems. By learning reward machines directly from demonstrations, agents can adapt to tasks without requiring manually specified task representations, making this approach scalable and practical for real-world applications, such as robotic manipulation and autonomous vehicle navigation [8, 9, 10, 11, 12]. This capability is especially valuable in environments where only high-level task-relevant features (aka, atomic propositions) are observable, underscoring the importance of learning RMs in advancing autonomous decision-making systems. Beyond autonomy applications, IRL and RM learning can also be used to infer the agent’s (e.g., humans’) intentions to design incentives or better decision making environments [13].
@@
Some of previous work on learning reward machines from data either assumes that the machine’s states are observed [14] or the reward is observed [10, 9, 15]. In the latter case the problem becomes finding a reward machine consistent with the observed input-output traces. Other work [16, 17, 18] infers reward machines by combining automata synthesis with reinforcement learning and querying the environment for experiences. Others [19, 20] use the standard $L^{*}$ algorithm for automata learning [21] to learn a consistent reward machine. This assumes access to an oracle that can answer membership and conjectures queries. There are also works that only use observations of atomic propositions [12], similar to us, however they are limited to single-stage goal reaching tasks, where the RM has a simple structure that is merely used to obtain dense rewards. In parallel, several works aim to infer an LTL specification from demonstrations satisfying and/or violating the specification [22, 23], requiring a potentially large, labeled data set. Since LTL learning problem is inherently ill-posed, several regularization techniques are used such as formula templates or concept classes.
@@
To the best of the authors’ knowledge, no prior work has formalized and solved the problem of learning reward machines from partially observed optimal policies directly without the need to observe the rewards or the machine’s state. The two main challenges of this setting are 1) partial observability (the reward is not observed, only the atomic propositions are observed), 2) partial reachability (not all transitions of the reward machine are visited in a given environment). In this work, we address these challenges by first characterizing what can be learned in this setting (i.e., an equivalence class of reward machines) and then proposing a SAT-based algorithm, which provably learns a reward machine equivalent to the underlying true one. The key insight of our algorithm is to identify pairs of atomic proposition prefixes, namely negative examples, that lead to different nodes of the underlying reward machine from the observable optimal prefix-based policy, and encoding these examples as constraints in the SAT problem. We demonstrate the efficacy of our algorithm with several examples.
@@
II Preliminaries and Problem Statement
II-A Markov Decision Processes and Reward Machines
@@
$\mathcal{T}_{s}(\mathcal{M})=\{(s_{0},\ldots,s_{t},\ldots)\in$ $(\mathcal{S})^{\omega}\mid\exists a_{0},\ldots,a_{t-1},\ldots,$
$\mathcal{P}(s_{t+1}\mid s_{t},a_{t})>0,\forall t\}.$
@@
$\mathrm{Pref}(\mathcal{L})=\{w\in(\mathrm{AP})^{*}\mid\exists x\in\mathcal{L},\,w\text{ is a prefix of }x\}.$ (1)
@@
$\mathrm{Reach}(\sigma)$ =$=\{s\in\mathcal{S}\mid\tau\in\mathcal{T}_{s}^{\mathrm{fin}}\text{s.t. }L(\tau)=\sigma,\tau_{end}=s\}.$
@@
As a running example, we borrow the patrol task from [2]. Consider the room grid world shown in Figure 1(a). It is a 4 by 4 grid where the agent can move in the four cardinal directions, with a small probability of slipping to neighboring cells. We color-code different cells to denote the proposition label associated with the corresponding cell. For example, all cells colored green have the high level proposition $\mathrm{A}$. The agent is tasked to patrol the rooms in the order $\mathrm{A}\to\mathrm{B}\to\mathrm{C}\to\mathrm{D}$. This is captured by the reward machine shown in Figure 1(b). Assume the agent’s state trajectory starts with $\tau=(\mathbf{a}_{1},\mathbf{a}_{2},\mathbf{a}_{3},\mathbf{b}_{4},\mathbf{c}_{1})$. The proposition sequence associated with $\tau$ is $\sigma=\mathrm{AAABC}$. The RM nodes traversed by following $\tau$ are $(u_{0},u_{1},u_{1},u_{1},u_{2},u_{3})$. Since $\mathbf{c}_{1}$ is the only state than can be reached with $\sigma$, we have that $\mathrm{Reach}(\sigma)=\{\mathbf{c}_{1}\}$. Similarly, $\sigma\in\mathrm{Paths}(u_{3})$.
@@
Figure 1: (a) The Room Grid World and (b) the Patrol Reward Machine.
II-B Markov Decision Process with a Reward Machine
A Markov decision process with a reward machine (MDP-RM) is a tuple $\mathcal{R}_{\mathcal{M}}=(\mathcal{M}/r,\mathcal{R},L)$ where $\mathcal{M}$ and $\mathcal{R}$ are defined as in Section II-A, and $L$ is a labeling function $L:\mathcal{S}\to\mathrm{AP}$.
@@ II-B Markov Decision Process with a Reward Machine
An MDP-RM can be equivalently seen as a product MDP $\mathcal{M}_{\mathrm{Prod}}=(\mathcal{S}^{\prime},\mathcal{A}^{\prime},\mathcal{P}^{\prime},\mu_{0}^{\prime},\gamma^{\prime},r^{\prime})$ where $\mathcal{S}^{\prime}=\mathcal{S}\times\mathcal{U}$, $\mathcal{A}^{\prime}=\mathcal{A}$, $\mathcal{P}^{\prime}(s^{\prime},u^{\prime}|s,u,a)=\mathcal{P}(s^{\prime}|s,a)\mathbf{1}(u^{\prime}=\delta_{\textbf{u}}(u,L(s^{\prime})))$, $\gamma^{\prime}=\gamma$, $\mu_{0}^{\prime}\in\Delta(\mathcal{S}\times\mathcal{U})$ with $\mu_{0}^{\prime}(s,u)=\mu_{0}(s)\mathbf{1}(u=u_{I})$ and $r^{\prime}(s,u,a,s^{\prime},u^{\prime})=\delta_{\mathbf{r}}(u,L(s^{\prime}))$. To make the notation compact, we denote the product state by $\bar{s}$, where $\bar{s}=(s,u)$. The product of an MDP model with a RM model is a product MDP model $\mathcal{G}\times\mathcal{M}_{L}=(\mathcal{S}^{\prime},\mathcal{A}^{\prime},\mathcal{T}^{\prime},\mu_{0}^{\prime},\gamma^{\prime})$ defined similarly.
@@
A trajectory of the product MDP $\mathcal{M}_{\mathrm{Prod}}$ is a sequence $(\bar{s}_{\emptyset},a_{\emptyset},\bar{s}_{0},a_{0},\bar{s}_{1},a_{1},\cdots)$, where $\bar{s}_{\emptyset}=(\emptyset,u_{I})$ and $a_{\emptyset}=\emptyset$. An initial state $s_{0}$ is sampled from $\mu_{0}$. The augmentation of $\bar{s}_{\emptyset}$ and $a_{\emptyset}$ at the start of the trajectory is to ensure that $s_{0}$ induces a transition in the reward machine. The reward machine thus transitions to $u_{0}=\delta_{\textbf{u}}(u_{I},L(s_{0}))$. The agent then takes action $a_{0}$ and transitions to $s_{1}$. Similarly, the reward machine transitions to $u_{1}=\delta_{\textbf{u}}(u_{0},L(s_{1}))$. The same procedure continues infinitely. We consider the product policy $\pi_{\mathrm{Prod}}:\mathrm{Dom_{Prod}}\to\Delta(\mathcal{A})$ where $\mathrm{Dom_{Prod}}\subseteq\mathcal{S}\times\mathcal{U}$ is the set of accessible $(s,u)$ pairs in the product MDP. This policy is a function that describes an agent’s behavior by specifying an action distribution at each state. We consider the Maximum Entropy Reinforcement Learning (MaxEntRL) objective given by:
@@
where $\lambda>0$ is a regularization parameter, and $\mathcal{H}(\pi(.|\bar{s}))=-\sum\limits_{a\in\mathcal{A}}\pi(a|\bar{s})\log(\pi(a|\bar{s}))$ is the entropy of the policy $\pi$. The expectation is with respect to the probability distribution $\mathbb{P}^{\pi}_{\mu_{0}}$, the induced distribution over infinite trajectories following $\pi$, $\mu_{0}$, and the Markovian transition kernel $\mathcal{P}^{\prime}$ [24]. The optimal policy $\pi_{\mathrm{Prod}}^{*}$, corresponding to a reward function $r^{\prime}$, is the maximizer of (2), i.e.,
@@
Optimal product MDP trajectories are trajectories of the product MDP generated using $\pi_{\mathrm{Prod}}^{*}$. We overload this definition to optimal trajectories of the MDP, which is generated from the optimal product MDP trajectories by simply removing the $u$ states. For the rest of the paper, optimal trajectories refer to the optimal trajectories of the MDP.
II-C Prefix Tree Policy
@@
where $\sigma\in\mathrm{Pref}(\mathcal{L}(\mathcal{M}_{L}))$. In particular, we say that the product policy $\pi_{\mathrm{Prod}}$ induces $\pi_{\mathrm{PTP}}$. We define the depth-$l$ restriction of a PTP as its restriction to the set $\left(\cup_{j=1}^{k}\mathcal{S}\times(\mathrm{AP})^{j}\right)\cap\mathrm{Dom_{PTP}}$. That is the policy associated to words $\sigma$ of length up to $l$. It is denoted by $\pi_{\mathrm{PTP}}^{l}$.
Remark 1
In this work, we assume that we have access to the depth-$l$ restriction of the PTP for a sufficiently large $l$ (the value of $l$ is discussed in Section III-C). This is a reasonable assumption because this quantity can be approximated arbitrary well by a finite set of optimal trajectory prefixes. Note that this is because the optimal trajectories of the MDP for the max entropy objective cover $\mathrm{Dom_{PTP}}$. However, the impact of such an approximation and related sample complexity questions are left for future work.
@@
Definition 1
Two reward machines are policy-equivalent with respect to a labeled MDP if the product policies obtained by solving problem (3) for each of the reward machines induce the same prefix tree policy defined as in Eq. (4). Among all the reward machines that are policy equivalent with respect to a labeled MDP, we define a minimal reward machine as one with the fewest number of nodes.
Several equivalence relations among reward machines in the literature are special cases of policy equivalence. For instance, when learning finite state machines from observed rewards [10, 25], two reward machines are said to be input-output equivalent if they produce the same reward sequence for the same atomic proposition sequence. Such input-output equivalent reward machines are clearly policy-equivalent. Furthermore, for a trivial, i.e., one state, reward machine, our definition reduces to the policy-equivalence definition in the standard inverse reinforcement learning problem [26, 27].
II-D Problem Statement

### v2

# Learning Reward Machines from Partially Observed Policies

@@ Abstract
Inverse reinforcement learning is the problem of inferring a reward function from an optimal policy or demonstrations by an expert. In this work, it is assumed that the reward is expressed as a reward machine whose transitions depend on atomic propositions associated with the state of a Markov Decision Process (MDP). Our goal is to identify the true reward machine using finite information. To this end, we first introduce the notion of a prefix tree policy which associates a distribution of actions to each state of the MDP and each attainable finite sequence of atomic propositions. Then, we characterize an equivalence class of reward machines that can be identified given the prefix tree policy. Finally, we propose a SAT-based algorithm that uses information extracted from the prefix tree policy to solve for a reward machine. It is proved that if the prefix tree policy is known up to a sufficient (but finite) depth, our algorithm recovers the exact reward machine up to the equivalence class. This sufficient depth is derived as a function of the number of MDP states and (an upper bound on) the number of states of the reward machine. These results are further extended to the case where we only have access to demonstrations from an optimal policy. Several examples, including discrete grid and block worlds, a continuous state-space robotic arm, and real data from experiments with mice, are used to demonstrate the effectiveness and generality of the approach.
1 Introduction
@@
Several frameworks exist for solving complex multi-staged tasks, including hierarchical reinforcement learning (HRL) (Pateria et al., 2021), reward machines (RMs) (Icarte et al., 2018) and linear temporal logic (LTL) specifications (Chou et al., 2020; Vaezipoor et al., 2021). HRL leverages a decomposition of tasks into subtasks, enabling agents to focus on solving smaller, manageable problems before integrating solutions into a higher-level policy (Sutton et al., 1999). On the other hand, RM and its generalizations (Corazza et al., 2022) encode task-specific knowledge as finite-state machines, capturing temporal dependencies and logical constraints in a concise and interpretable manner, similar to LTL. This structure simplifies policy learning and improves efficiency, especially in environments with long horizons or sparse rewards.
@@
As an extension to inverse reinforcement learning (IRL) (Ng et al., 2000), one could ask the question of learning RMs, which enables agents to autonomously extract structured representations of tasks, significantly enhancing their ability to solve complex, temporally extended problems. By learning reward machines directly from demonstrations, agents can adapt to tasks without requiring manually specified task representations, making this approach scalable and practical for real-world applications, such as robotic manipulation and autonomous vehicle navigation (Camacho et al., 2019; Icarte et al., 2023; Xu et al., 2020; Baert et al., 2024; Camacho et al., 2021). This capability is especially valuable in environments where high-level task-relevant features (aka, atomic propositions) are observable, underscoring the importance of learning RMs in advancing autonomous decision-making systems. Beyond autonomy applications, IRL and RM learning can also be used to infer the agent’s (e.g., humans’) intentions to design incentives or better decision making environments (Nitschke et al., 2024).
@@
Some of previous work on learning reward machines from data either assumes that the machine’s states are observed (Araki et al., 2019) or the reward is observed (Xu et al., 2020; Icarte et al., 2023; Hu et al., 2024; Abate et al., 2023). In the latter case the problem becomes finding a reward machine consistent with the observed input-output traces. Other work (Hasanbeig et al., 2024; 2021; Furelos-Blanco et al., 2020) infers reward machines by combining automata synthesis with reinforcement learning and querying the environment for experiences. Others (Xu et al., 2021; Memarian et al., 2020) use the standard $L^{*}$ algorithm for automata learning (Angluin, 1987) to learn a consistent reward machine. This assumes access to an oracle that can answer membership and conjectures queries. There are also works that only use observations of atomic propositions (Camacho et al., 2021; Baert et al., 2024), similar to us; however, they are limited to single-stage goal reaching tasks, where the RM has a simple structure that is merely used to obtain dense rewards. In parallel, several works aim to infer an LTL specification from demonstrations satisfying and/or violating the specification (Neider and Gavran, 2018; Vazquez-Chanlatte and Seshia, 2020), requiring a potentially large, labeled data set. Since LTL learning problem is inherently ill-posed, several regularization techniques are used such as formula templates or concept classes.
@@
To the best of the authors’ knowledge, no prior work has formalized and solved the problem of learning reward machines from partially observed optimal policies directly without the need to observe the rewards or the machine’s state. The two main challenges of this setting are 1) partial observability (the reward is not observed, only the atomic propositions are observed), 2) partial reachability (not all transitions of the reward machine are visited in a given environment). In this work, we address these challenges by first characterizing what can be learned in this setting (i.e., an equivalence class of reward machines) and then proposing a SAT-based algorithm, which provably learns a reward machine equivalent to the underlying true one. The key insight of our algorithm is to identify pairs of atomic proposition prefixes, namely negative examples, that lead to different nodes of the underlying reward machine from the observable optimal prefix-based policy, and encoding these examples as constraints in the SAT problem. We show that our method can be applied even when the optimal policy is accessible only through a finite set of optimal trajectories. To this end, we approximate the policy from the data and replace the SAT problem with a variant called weighted MAX-SAT that provides robustness to incorrectly labeled negative examples. We demonstrate the efficacy of our algorithm in diverse settings, including grid-based MDPs, a robotic control task, and a real-world dataset of mouse navigation.
@@
2 Preliminaries and Problem Statement
2.1 Markov Decision Processes and Reward Machines
@@
$\mathcal{T}_{s}(\mathcal{M})=\{(s_{0},s_{1},\ldots)\in$ $\mathcal{S}^{\omega}\mid\exists(a_{0},a_{1},\ldots)\in\mathcal{A}^{\omega}:\mathcal{P}(s_{t+1}\mid s_{t},a_{t})>0,\forall t\}.$
@@
$\mathrm{Pref}(\mathcal{L})=\{w\in(\mathrm{AP})^{*}\mid\exists x\in\mathcal{L},\text{ s.t. }w\text{ is a prefix of }x\}.$
@@
$\mathrm{Reach}(\sigma)$ =$=\{s\in\mathcal{S}\mid\tau\in\mathcal{T}_{s}^{\mathrm{fin}}\text{s.t. }L(\tau)=\sigma,\tau_{|\tau|}=s\}.$
@@
As a running example, we borrow the patrol task from (Icarte et al., 2018). Consider the room grid world shown in Figure 1(a). It is a 4 by 4 grid where the agent can move in the four cardinal directions, with a small probability of slipping to neighboring cells. We color-code different cells to denote the proposition label associated with the corresponding cell. For example, all cells colored green have the high level proposition $\mathrm{A}$. The agent is tasked to patrol the rooms in the order $\mathrm{A}\to\mathrm{B}\to\mathrm{C}\to\mathrm{D}$. This is captured by the reward machine shown in Figure 1(b). Assume the agent’s state trajectory starts with $\tau=(\mathbf{a}_{1},\mathbf{a}_{2},\mathbf{a}_{3},\mathbf{b}_{4},\mathbf{c}_{1})$. The proposition sequence associated with $\tau$ is $\sigma=\mathrm{AAABC}$. The RM nodes traversed by following $\tau$ are $(u_{0},u_{1},u_{1},u_{1},u_{2},u_{3})$. Since $\mathbf{c}_{1}$ is the only state than can be reached with $\sigma$, we have that $\mathrm{Reach}(\sigma)=\{\mathbf{c}_{1}\}$. Similarly, $\sigma\in\mathrm{Paths}(u_{3})$. Each transition in the reward machine gives a reward of zero, except the transition from $u_{3}$ to $u_{0}$, i.e., $\delta_{\mathbf{r}}(u,l)=1$ if $(u,l)=(u_{3},D)$, and zero otherwise.
@@
Figure 1: (a) The room grid world. (b) The patrol reward machine. (c) The room grid world with a hallway.
2.2 Markov Decision Process with a Reward Machine
@@
A Markov decision process with a reward machine (MDP-RM) is a tuple $\mathcal{R}_{\mathcal{M}}=(\mathcal{M}/r,\mathcal{R},L)$ where $\mathcal{M}$ and $\mathcal{R}$ are defined as in Section 2.1, and $L$ is a labeling function $L:\mathcal{S}\to\mathrm{AP}$. An MDP-RM can be equivalently seen as a product MDP $\mathcal{M}_{\mathrm{Prod}}=(\mathcal{S}^{\prime},\mathcal{A}^{\prime},\mathcal{P}^{\prime},\mu_{0}^{\prime},\gamma^{\prime},r^{\prime})$ where $\mathcal{S}^{\prime}=\mathcal{S}\times\mathcal{U}$, $\mathcal{A}^{\prime}=\mathcal{A}$, $\mathcal{P}^{\prime}(s^{\prime},u^{\prime}|s,u,a)=\mathcal{P}(s^{\prime}|s,a)\mathbf{1}(u^{\prime}=\delta_{\textbf{u}}(u,L(s^{\prime})))$, $\gamma^{\prime}=\gamma$, $\mu_{0}^{\prime}\in\Delta(\mathcal{S}\times\mathcal{U})$ with $\mu_{0}^{\prime}(s,u)=\mu_{0}(s)\mathbf{1}(u=u_{I})$ and $r^{\prime}(s,u,a,s^{\prime},u^{\prime})=\delta_{\mathbf{r}}(u,L(s^{\prime}))$. To make the notation compact, we denote the product state by $\bar{s}=(s,u)$. The product of an MDP model with a RM model is a product MDP model $\mathcal{G}\times\mathcal{M}_{L}=(\mathcal{S}^{\prime},\mathcal{A}^{\prime},\mathcal{T}^{\prime},\mu_{0}^{\prime},\gamma^{\prime})$ defined similarly.
@@
A trajectory of the product MDP $\mathcal{M}_{\mathrm{Prod}}$ is a sequence $(\bar{s}_{\emptyset},a_{\emptyset},\bar{s}_{0},a_{0},\bar{s}_{1},a_{1},\cdots)$, where $\bar{s}_{\emptyset}=(\emptyset,u_{I})$ and $a_{\emptyset}=\emptyset$. An initial state $s_{0}$ is sampled from $\mu_{0}$. The introduction of $\bar{s}_{\emptyset}$ and $a_{\emptyset}$ at the start of the trajectory is to ensure that $s_{0}$ induces a transition in the reward machine. The reward machine thus transitions to $u_{0}=\delta_{\textbf{u}}(u_{I},L(s_{0}))$. The agent then takes action $a_{0}$ and transitions to $s_{1}$. Similarly, the reward machine transitions to $u_{1}=\delta_{\textbf{u}}(u_{0},L(s_{1}))$. The same procedure continues infinitely. We consider the product policy $\pi_{\mathrm{Prod}}:\mathrm{Dom_{Prod}}\to\Delta(\mathcal{A})$ where $\mathrm{Dom_{Prod}}\subseteq\mathcal{S}\times\mathcal{U}$ is the set of accessible $(s,u)$ pairs in the product MDP. This policy is a function that describes an agent’s behavior by specifying an action distribution at each state. We consider the Maximum Entropy Reinforcement Learning (MaxEntRL) objective given by:
@@
where $\lambda>0$ is a regularization parameter, and $\mathcal{H}(\pi(.|\bar{s}))=-\sum\limits_{a\in\mathcal{A}}\pi(a|\bar{s})\log(\pi(a|\bar{s}))$ is the entropy of the policy $\pi$. The expectation is with respect to the probability distribution $\mathbb{P}^{\pi}_{\mu_{0}}$, the induced distribution over infinite trajectories following $\pi$, $\mu_{0}$, and the Markovian transition kernel $\mathcal{P}^{\prime}$ (Ziebart et al., 2008). The optimal policy $\pi_{\mathrm{Prod}}^{*}$, corresponding to a reward function $r^{\prime}$, is the maximizer of (1), i.e.,
@@
Optimal product MDP trajectories are trajectories of the product MDP generated using $\pi_{\mathrm{Prod}}^{*}$. We overload this definition to optimal trajectories of the MDP, which is generated from the optimal product MDP trajectories by simply removing the $u$ states. For the rest of the paper, optimal trajectories or demonstrations refer to the optimal trajectories of the MDP.
2.3 Prefix Tree Policy
@@
where $\sigma\in\mathrm{Pref}(\mathcal{L}(\mathcal{M}_{L}))$. In particular, we say that the product policy $\pi_{\mathrm{Prod}}$ induces $\pi_{\mathrm{PTP}}$. We define the depth-$l$ restriction of a PTP as its restriction to the set $\left(\cup_{j=1}^{l}\mathcal{S}\times(\mathrm{AP})^{j}\right)\cap\mathrm{Dom_{PTP}}$. That is the policy associated to words $\sigma$ of length up to $l$. It is denoted by $\pi_{\mathrm{PTP}}^{l}$.
@@
Definition 1.
Two reward machines are policy-equivalent with respect to a labeled MDP model if the product policies obtained by solving problem (2) for each of the reward machines induce the same prefix tree policy defined as in (3). Among all the reward machines that are policy equivalent with respect to a labeled MDP, we define a minimal reward machine as one with the fewest number of nodes.
Several equivalence relations among reward machines in the literature are special cases of policy equivalence. For instance, when learning finite state machines from observed rewards (Xu et al., 2020; Giantamidis et al., 2021), two reward machines are said to be input-output equivalent if they produce the same reward sequence for the same atomic proposition sequence. Such input-output equivalent reward machines are clearly policy-equivalent. Furthermore, for a trivial, i.e., one state, reward machine, our definition reduces to the policy-equivalence definition in the standard inverse reinforcement learning problem (Shehab et al., 2024; Cao et al., 2021).
2.4 Problem Statement